Ellen wants to make a walking path around the edge of her property. Her property is 500ft x 400ft. The walking path will be 3 ft wide. How much area in square feet will the walking path take up oon her property?
Answer by ankor@dixie-net.com(22740)

Ellen wants to make a walking path around the edge of her property.
Her property is 500ft x 400ft. The walking path will be 3 ft wide.
How much area in square feet will the walking path take up on her property?
:
Find the area of the property: 500 * 400 = 200000 sq/ft
:
The area inside the path:
dimensions: (500-6) by (400-6) = 494 * 394 = 194636
then
200000 - 194636 = 5,364 sq/ft taken by the path